Find an equation for the line with y-intercept 3 that is perpendicular to the line y=2/3x-4.
Answer by malakumar_kos@yahoo.com(315)


Find an equation for the line with y-intercept 3 that is perpendicular to the line y=2/3x-4.
Let the required eq'n be y = mx+c
if the y- intercept is 3 then c= 3
To find the slope , the two lines are perpendicular hence product of their
slopes = -1. In the given eq'n slope M = 2/3

slope of the required line = m = ?
m*M = -1 or m = -1/M = -1/2/3 = -3/2
substituting the values of m and c we get
y = (-3/2)x+3 is the solution.